OpenStudy (anonymous):

Solve for variable exactly with all solutions: x(2x+3)=44

OpenStudy (hawkfalcon):

It's a quadratic equation.

OpenStudy (homeworksucks):

2x^2+3x=44 2x^2+3x-44=0 quadratic equation formula-\[x=(-b \pm( \sqrt{b ^{2}-4ac}))\div2a\] plugging it in, we get two solutions- -5.5 and 4. Hope this helps

OpenStudy (campbell_st):

distribute and write everything on the left hand side of the equals sign 2x^2 + 3x - 44 = 0 this can be factorises or solved using the general quadratic formula \[x = \frac{ -b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = 2, b = 3 and c = -44